OpenStudy (anonymous):

hockey teams recieve 2 points when they win and 1 point when they tie. One season a team won a championship with 54 points. They won 12 more games than they tied.How many wins and ties did the team have?

OpenStudy (amistre64):

w = 2 t = 1 54 = w + t w = t +12 ?

OpenStudy (anonymous):

54=2(12+x)+x?

OpenStudy (amistre64):

sats looks good

OpenStudy (anonymous):

54 = 2x + y x = 12 + y therefore, 54 = 2(12 +y) + y = 24 +3y y = 10, which means that x = 22 22 wins, 10 ties

OpenStudy (anonymous):

whew. 54=24+3x 3=10 i am a little slow today.

OpenStudy (anonymous):

meant x = 10

OpenStudy (anonymous):

that's ok so am I

OpenStudy (anonymous):

amistre64 i think i figured out the \[\frac{1}{1-2x}\] thing maybe.